# Maintenance Optimization for Asset Networks with Unknown Degradation Parameters

### v2

@@ 3.5 Structural properties of the underlying Markov decision process
In this section, we establish several structural properties of optimal replacement policies $\pi^{\textbf{L}_{2}}_{*}$ of the underlying MDP of the POMDP model presented in Section 3.3.
@@
From Theorem 1, it follows that optimal policies are state-dependent threshold policies. Figure 1 shows an example of an optimal policy for a 2-asset instance retrieved via policy iteration. Note that because of the shared setup costs, the optimal action in the set of states $\{(12,x_{2})\mid x_{2}\in\{15,\ldots,19\}\}\cup\{(13,x_{2})\mid x_{2}\in\{14,\ldots,19\}\}\cup\{(14,x_{2})\mid x_{2}\in\{13,\ldots,19\}\}\cup\{(x_{1},12)\mid x_{1}\in\{15,\ldots,19\}\}\cup\{(x_{1},13)\mid x_{1}\in\{14,\ldots,19\}\}\cup\{(x_{1},14)\mid x_{1}\in\{13,\ldots,19\}\}$ is to “maintain all”, even though the optimal PM threshold appears to be $15$. This example demonstrates that generally computing optimal policies is not straightforward.
Figure 1: An optimal policy $\pi^{\textbf{L}_{2}}_{*}$ of the underlying MDP for $\Lambda_{m}\overset{d}{=}5$ and $Y^{(i)}_{m}\overset{d}{=}1$.
@@
Let $X$ and $Y$ be two random variables such that $\mathbb{P}(X>x)\leq\mathbb{P}(Y>x)$ for all $x\in\mathbb{R}$. Then $X$ is said to be smaller than $Y$ in the usual stochastic order, denoted by $X\leq_{\textrm{st}}Y$ (Shaked and Shanthikumar, 2007, 1.A.1).
@@
Let $x=(x_{1},\lambda_{1},\phi_{1},\ldots,x_{M},\lambda_{M},\phi_{M})$ denote a state. Let $\mathcal{M}(x)\subseteq\mathcal{M}$ again the set of machines for which it is optimal to do maintenance in state $x$. For any $x^{\prime}=(x_{1},\lambda^{\prime}_{1},\phi^{\prime}_{1},\ldots,x_{M},\lambda^{\prime}_{M},\phi^{\prime}_{M})\in\mathcal{S}$ that represents a state with worse degradation parameters than $x$, i.e.,
@@
Theorem 2 essentially states that if a state $x^{\prime}$ is “worse” than another state $x$ in terms of degradation parameters, then the optimal maintenance action remains the same. Specifically, if the degradation parameters $\lambda_{m}$ for a machine are greater or equal in state $x^{\prime}$ compared to state $x$, and the shock size distributions $Y^{\prime}_{m}$ are larger than $Y_{m}$ in the usual stochastic order, then the same maintenance actions should be applied in both states. Note that Theorem 2 applies to a broad range of MDPs with shock size distributions from the one-parameter exponential family, particularly the geometric distribution supported on $\mathbb{Z}_{+}$.
While solving for an optimal policy explicitly can be computationally infeasible for larger asset networks, the derived structural properties provide valuable insights that can guide policy design. In DRL, heuristic strategies play a crucial role: they serve not only as benchmarks for evaluating trained agents but also as initialization strategies that help kickstart learning in challenging environments. Verleijsdonk et al. (2024), for example, use carefully designed heuristic policies to initialize the API algorithm and demonstrate that such initialization significantly reduces the number of iterations needed to reach a near-optimal solution.
@@ 4.1 Two-threshold control limit heuristic
The two-threshold control limit heuristic is an offline heuristic approach requiring at least information level $\textbf{L}_{1}$. The two-threshold control limit involves two distinct thresholds: one for initiating preventive maintenance (PM) and another for opportunistic preventive maintenance (OPM). The PM threshold $\tau_{m}^{\textrm{PM}}\in(0,\xi_{m}]$ triggers scheduled maintenance activities when the $m$-th asset’s degradation signal is greater than or equal to $\tau_{m}^{\textrm{PM}}$. Meanwhile, the OPM threshold $\tau_{m}^{\textrm{OPM}}\in(0,\tau_{m}^{\textrm{PM}}]$ is set at a lower degradation level, prompting maintenance activities when another asset is already undergoing maintenance. This approach capitalizes on the opportunity to address potential issues while sharing maintenance setup costs, thereby reducing the overall costs. We refer to this approach as $\pi^{\textbf{L}_{1}}_{\mathcal{N}}$ because it is the most naive heuristic approach; it does not explicitly consider component heterogeneity, yet it is widely used in practice. Note that setting $\tau_{m}^{\textrm{OPM}}=\tau_{m}^{\textrm{PM}}\geq\xi_{m}$ for all $m\in\mathcal{M}$ yields a reactive heuristic, which we shall denote by $\pi^{\textbf{L}_{1}}_{\mathcal{R}}$.
@@ 4.2 Integrated Bayes heuristic
The integrated Bayes heuristic approach is adopted from Drent et al. (2023, Section 5.3), and applies a state-space truncation to solve the objective of Eq. (8) in the single asset scenario using value iteration. We extend the policy to the multi-asset scenario by applying the obtained policy $\pi^{\textbf{L}_{1}}_{\mathcal{I}}$ to each asset individually, thus ignoring the economic dependence. Due to the curse of dimensionality, solving the objective of Eq. (8) for larger asset networks is computationally intractable.
@@ 4.3 Approximate policy iteration for Bayesian Markov decision processes
To learn maintenance policies for economically coupled assets in the context of BMDPs, we employ a variation of approximate policy iteration. Specifically, we utilize deep controlled learning (DCL), which was first introduced by Temizöz et al. (2025). Verleijsdonk et al. (2024) have shown that API/DCL can effectively learn maintenance (together with the dispatching strategies) for an industrial-scale network of homogeneous assets. Moreover, DCL has demonstrated superior performance in inventory problems compared to other DRL algorithms including proximal policy optimization and asynchronous advantage actor-critic (Temizöz et al., 2025).
To apply DCL to BMDPs, we introduce various novel techniques: (i) randomized action selection to avoid indexation bias, (ii) sampling from the prior distributions to directly train neural network policies for BMDPs, (iii) leverage heuristic information to limit the action space in a significant number of states, and (iv) an open-loop feedback feature vector that enables the application of neural network policies trained in an $\textbf{L}_{2}$ setting to be applied in an $\textbf{L}_{1}$ setting.
@@ 4.3.1 Deep controlled learning for Bayesian Markov decision processes
Following the approach of Verleijsdonk et al. (2024), we apply DCL to train a neural network for sequential action selection in asset management, thereby significantly reducing the complexity of the action space. To avoid indexation bias, we randomly select a new permutation $\sigma\in S(\mathcal{M})$ at each decision epoch, ensuring that the order in which actions are chosen for machines is randomized. Following each action selection, the input state is updated with the effects of the action before deciding on an action for the next asset. More specifically, the state transition $h\rightarrow h^{\prime}$ is decomposed into two stages: The first stage is governed by the deterministic outcomes of the selected actions $a\in\mathcal{U}(h)$, while the second stage is influenced by the random progression of the degradation processes. In more detail, $h\rightarrow h^{\prime}$ is decomposed into $h\overset{a_{\sigma{(1)}}}{\rightarrow}h^{a_{\sigma{(1)}}}\overset{a_{\sigma{(2)}}}{\rightarrow}\ldots\overset{a_{\sigma{(M)}}}{\rightarrow}h^{a_{\sigma{(M)}}}=:h^{a}$ and to $h^{a}\overset{t\rightarrow t+1}{\longrightarrow}h^{\prime}$. The sequence in which the actions $a_{1},\ldots,a_{M}$ are handled does not matter, so we will only describe the processing of the action for the $m$-th asset.
@@
We now provide a concise overview of the application of DCL to BMDPs (on the application of DCL to MDPs, we refer to Verleijsdonk et al. (2024) and to Temizöz et al. (2025)). Under the information level L, the DCL algorithm comprises the following three steps:
@@
    Construct a data set $\mathcal{D}$ using $\pi^{\textbf{L}}_{0}$ containing state-action mappings.
@@ 4.3.2 Initial policies and feature representations
Verleijsdonk et al. (2024, Section 5.3) argue that initiating DCL with an appropriate policy $\pi^{\textbf{L}}_{0}$ considerably reduces computation times. The optimized two-threshold control limit heuristic satisfies all relevant listed properties since it explores sufficiently many states and has low computational complexity (as opposed to the integrated Bayes heuristic from Section 4.2). Moreover, the optimized PM and OPM thresholds $\tau_{m}^{\text{PM}}$ and $\tau_{m}^{\text{OPM}}$ can be leveraged to effectively restrict the individual, state-dependent action sets $\mathcal{U}_{m}(h)$ for a significant number of states. This is particularly useful when the failure threshold $\xi_{m}$ is relatively large. Specifically, we train and evaluate the policy improvements of the two-threshold control limit on a restricted action space defined as $\tilde{\mathcal{U}}(h)=\tilde{\mathcal{U}}_{1}(h)\times\ldots\times\tilde{\mathcal{U}}_{M}(h)$, where for each $m\in\mathcal{M}$, the individual action set $\tilde{\mathcal{U}}_{m}(h)$ is given by:
@@
Here, for each $m\in\mathcal{M}$, $\delta_{m}\in[0,1]$ and $\zeta_{m}\geq 1$ should be chosen conservatively. For all numerical experiments in this work, we set $\delta_{m}\equiv 0.5$ and $\zeta_{m}\equiv 1.5$.
The feature representation of a state depends on the information available to the manager. We propose a feature design that communicates the state information for each asset in a manner that is customized to the specific asset for which we are currently determining an action. Verleijsdonk et al. (2024, Section 7.1) demonstrate that using such a feature design results in less training variability and consistently produces policies that perform significantly better.
@@
that is, the feature vector $f^{\textbf{L}_{2}}_{1}(h)$ includes an information block $(x_{m},\lambda_{m},p_{m},\iota_{m})$ for each $m\in\mathcal{M}$ that can be derived from $h$, along with one additional feature. Here, $x_{m}$ is the observed degradation level of the $m$-th asset, and the entries $\lambda_{m}$ and $p_{m}$ denote its degradation parameters. The last block entry $\iota_{m}$ indicates whether we are currently selecting an action for asset $m$. Lastly, the additional feature $\eta$ indicates whether there is an opportunity for OPM. More specifically, $\eta$ equals $1$ if maintenance on at least one asset is mandatory or if the maintenance action has already been selected in the current decision epoch, and 0 otherwise. Note that the dimension of the feature vector is $r=4M+1$.
@@
$(\lambda^{\textrm{Bayes}}_{m},p^{\textrm{Bayes}}_{m})=\left(\frac{\alpha_{m}(t)}{\beta_{m}(t)},\frac{a_{m}(t)}{a_{m}(t)+b_{m}(t)}\right),$ (9)
where $\alpha_{m}(t),\beta_{m}(t),a_{m}(t)$ and $b_{m}(t)$ are defined in Eqs. (4)–(7).
@@ 5 Simulation study
We compare the performance of each proposed solution approach with that of a policy optimized under information level $\textbf{L}_{2}$. We restrict our analysis to the case where $M=2$ as it facilitates visualization of the trained policies. Furthermore, we assume that all component replacements come from the same pool of components, meaning that $\Lambda_{m}\overset{\textrm{d}}{\equiv}\Lambda$ and $\Phi_{m}\overset{\textrm{d}}{\equiv}\Phi$. For each instance of the simulation study, the true hyperparameters of the gamma distribution $\Lambda$ and the beta distribution $\Phi$ that model the population heterogeneity are denoted by $\alpha$, $\beta$, $a$ and $b$, and are listed in Table 1. The means of the distributions $\Lambda$ and $\Phi$ are denoted by $\mu_{\Lambda}$ and $\mu_{\Phi}$, respectively. These instances are selected to reflect increasing volatility in component heterogeneity, represented by an increasing coefficient of variation (CV) of the distributions, as well as a range of representative cost parameters.
Instance $M$ $\xi_{m}$ $\mu_{\Lambda}$ $\textrm{CV}_{\Lambda}$ $\mu_{\Phi}$ $\textrm{CV}_{\Phi}$ $\alpha$ $\nicefrac{{1}}{{\beta}}$ $a$ $b$ $c_{m}^{\textrm{PM}}$ $c_{m}^{\textrm{CM}}$ $c^{\textrm{ST}}$ $\gamma$ I.1 22 2020 11 $0.3$ $\frac{1}{2}$ $0.01$ $11.11$ $0.09$ $4999.5$ $4999.5$ 1 5 1 0.99 I.2 22 2020 11 $0.6$ $\frac{1}{2}$ $0.02$ $2.78$ $0.36$ $1249.5$ $1249.5$ 1 10 1 0.99
Table 1: Hyperparameter settings and cost structures (adopted from Drent et al. (2023, Section 5)) considered in the simulation study.
All performance results in this section are retrieved using $10^{6}$ repetitions of length $10^{3}$ time units, if applicable. The reported half-widths represent asymptotic $95\%$ confidence intervals and account solely for variability within the model.

### v3

@@ 3.5 Structural properties of the underlying Markov decision process
In this section, we establish several structural properties of optimal replacement policies $\pi^{\textbf{L}_{2}}_{*}$ of the underlying MDP of the POMDP model presented in Section 3.3. We provide the proof of Theorem 1 and Theorem 2 explicitly for the case $M=2$ in A. The argument extends analogously to the case $M>2$.
@@
From Theorem 1, it follows that optimal policies are state-dependent threshold policies. Figure 2 shows an example of an optimal policy for a 2-asset instance retrieved via policy iteration. Note that, due to the presence of shared setup costs, the optimal action for all states within the black boundary is to “maintain all”, despite the fact that the optimal PM threshold for an asset considered in isolation is $15$. The solid black boundary is inclusive, whereas the dashed boundary is exclusive. This example highlights the complexity involved in determining optimal policies.
Figure 2: An optimal policy $\pi^{\textbf{L}_{2}}_{*}$ of the underlying MDP for the 2-asset instance with parameters $(c_{m}^{\textrm{PM}},c_{m}^{\textrm{CM}},c_{m}^{\textrm{ST}},\gamma)=(1,10,3,0.99)$, $\Lambda_{m}\overset{\textrm{d}}{=}5$ and $Y^{(i)}_{m}\overset{\textrm{d}}{=}1$.
@@
Let $X$ and $Y$ be two random variables such that $\mathbb{P}(X>x)\leq\mathbb{P}(Y>x)$ for all $x\in\mathbb{R}$. Then $X$ is said to be smaller than $Y$ in the usual stochastic order, denoted by $X\leq_{\textrm{st}}Y$ [Shaked and Shanthikumar, 2007, 1.A.1].
@@
Let $x=(x_{1},\lambda_{1},\phi_{1},\ldots,x_{M},\lambda_{M},\phi_{M})$ denote a state. Let $\mathcal{M}(x)\subseteq\mathcal{M}$ denote the set of machines for which it is optimal to do maintenance in state $x$. For any $x^{\prime}=(x_{1},\lambda^{\prime}_{1},\phi^{\prime}_{1},\ldots,x_{M},\lambda^{\prime}_{M},\phi^{\prime}_{M})$ that represents a state with worse degradation parameters than $x$, i.e.,
@@
Theorem 2essentially states that if a state $x^{\prime}$ is “worse” than another state $x$ in terms of degradation parameters, then the optimal maintenance action remains the same. Specifically, if the degradation parameters $\lambda^{\prime}_{m}$ in state $x^{\prime}$ are greater than or equal to $\lambda_{m}$ in state $x$, and the shock size distributions $Y^{\prime}_{m}$ are larger than $Y_{m}$ in the usual stochastic order, then the same maintenance actions should be applied in both states. Note that Theorem 2 applies to a broad range of MDPs with shock size distributions from the one-parameter exponential family, particularly the geometric distribution supported on $\mathbb{Z}_{+}$.
While solving for an optimal policy explicitly can be computationally infeasible for larger asset networks, the derived structural properties provide valuable insights that can guide policy design. Although trained DRL agents generally do not satisfy structural guarantees such as the monotonicity properties derived here, the results remain valuable in broader DRL contexts, as they provide theoretical guidance for designing effective heuristic policies and for reducing the complexity of the DRL algorithm by shrinking the action space. In DRL, such heuristic strategies play a crucial role: they serve not only as benchmarks for evaluating trained agents but also as initialization strategies that help kickstart learning in challenging environments. Verleijsdonk et al. [2024], for example, use carefully designed heuristic policies to initialize the API algorithm and demonstrate that such initialization significantly reduces the number of iterations needed to reach a near-optimal solution.
@@ 4.1 Two-threshold control limit heuristic
The two-threshold control limit heuristic is an offline heuristic approach requiring at least information level $\textbf{L}_{1}$. This heuristic involves two distinct thresholds: one for initiating preventive maintenance (PM) and another for opportunistic preventive maintenance (OPM). The PM threshold $\tau_{m}^{\textrm{PM}}\in(0,\xi_{m}]$ triggers scheduled maintenance activities when the $m$-th asset’s degradation signal is greater than or equal to $\tau_{m}^{\textrm{PM}}$. Meanwhile, the OPM threshold $\tau_{m}^{\textrm{OPM}}\in(0,\tau_{m}^{\textrm{PM}}]$ is set at a lower degradation level, prompting maintenance activities when another asset is already undergoing maintenance. This approach capitalizes on the opportunity to address potential issues while sharing maintenance setup costs, thereby reducing the overall costs. We refer to this approach as $\pi^{\textbf{L}_{1}}_{\mathcal{N}}$ because it is the most naive heuristic approach; it does not explicitly consider component heterogeneity, yet it is widely used in practice. Note that setting $\tau_{m}^{\textrm{OPM}}=\tau_{m}^{\textrm{PM}}\geq\xi_{m}$ for all $m\in\mathcal{M}$ yields a reactive heuristic, which we shall denote by $\pi^{\textbf{L}_{1}}_{\mathcal{R}}$.
@@ 4.2 Integrated Bayes heuristic
The integrated Bayes heuristic approach is adopted from Drent et al. [2023, Section 5.3], and applies a state-space truncation to solve the objective of Eq. (8) in the single asset scenario using value iteration. We extend the policy to the multi-asset scenario by applying the obtained policy $\pi^{\textbf{L}_{1}}_{\mathcal{I}}$ to each asset individually, thus ignoring the economic dependence. Due to the curse of dimensionality, solving the objective of Eq. (8) for larger asset networks is computationally intractable.
@@ 4.3 Approximate policy iteration for Bayesian Markov decision processes
To learn maintenance policies for economically coupled assets in the context of BMDPs, we employ a variation of approximate policy iteration. Specifically, we utilize deep controlled learning (DCL), which was first introduced by Temizöz et al. [2025]. Verleijsdonk et al. [2024] have shown that API/DCL can effectively learn maintenance policies (together with dispatching strategies) for an industrial-scale network of homogeneous assets, and that it outperforms several baseline heuristics, including decomposition-based methods and solutions derived from combinatorial optimization algorithms. Moreover, DCL has demonstrated superior performance in inventory problems compared to other DRL algorithms including proximal policy optimization and asynchronous advantage actor-critic [Temizöz et al., 2025].
To apply DCL to BMDPs, we introduce various novel techniques: (i) randomized action selection to avoid indexation bias, (ii) sampling from the prior distributions to directly train neural network policies for BMDPs, (iii) leveraging heuristic information to limit the action space in a significant number of states, and (iv) an open-loop feedback feature vector that enables the application of neural network policies trained in an $\textbf{L}_{2}$ setting to be applied in an $\textbf{L}_{1}$ setting.
@@ 4.3.1 Deep controlled learning for Bayesian Markov decision processes
Following the approach of Verleijsdonk et al. [2024], we apply DCL to train a neural network for sequential action selection in asset management, thereby significantly reducing the complexity of the action space. To avoid indexation bias, we randomly select a new permutation $\sigma\in S(\mathcal{M})$, where $S(\mathcal{M})$ denotes the set of all permutations of $\mathcal{M}$, at each decision epoch, ensuring that the order in which actions are chosen for machines is randomized. Following each action selection, the input state is updated with the effects of the action before the algorithm decides on an action for the next asset. More specifically, the state transition $h\rightarrow h^{\prime}$ is decomposed into two stages: The first stage is governed by the deterministic outcomes of the selected actions $a\in\mathcal{U}(h)$, while the second stage is influenced by the random progression of the degradation processes. In more detail, $h\rightarrow h^{\prime}$ is decomposed into $h\overset{a_{\sigma{(1)}}}{\rightarrow}h^{a_{\sigma{(1)}}}\overset{a_{\sigma{(2)}}}{\rightarrow}\ldots\overset{a_{\sigma{(M)}}}{\rightarrow}h^{a_{\sigma{(M)}}}=:h^{a}$ and to $h^{a}\overset{t\rightarrow t+1}{\longrightarrow}h^{\prime}$. The sequence in which the actions $a_{1},\ldots,a_{M}$ are handled does not matter, so we will only describe the processing of the action for the $m$-th asset.
@@
We now provide a concise overview of the application of DCL to BMDPs (on the application of DCL to MDPs, we refer to Verleijsdonk et al. [2024] and to Temizöz et al. [2025]). Under the information level L, the DCL algorithm comprises the following three steps:
@@
    Using $\pi^{\textbf{L}}_{0}$, construct a data set $\mathcal{D}$ containing state-action mappings.
@@ 4.3.2 Initial policies and feature representations
Verleijsdonk et al. [2024, Section 5.3] argue that initiating DCL with an appropriate policy $\pi^{\textbf{L}}_{0}$ considerably reduces computation times. The optimized two-threshold control limit heuristic satisfies all relevant listed properties since it explores sufficiently many states and has low computational complexity (as opposed to the integrated Bayes heuristic from Section 4.2). Moreover, the optimized PM and OPM thresholds $\tau_{m}^{\textrm{PM}}$ and $\tau_{m}^{\textrm{OPM}}$ can be leveraged to effectively restrict the individual, state-dependent action sets $\mathcal{U}_{m}(h)$ for a significant number of states. This is particularly useful when the failure threshold $\xi_{m}$ is relatively large. Specifically, we train and evaluate the policy improvements of the two-threshold control limit on a restricted action space defined as $\tilde{\mathcal{U}}(h)=\tilde{\mathcal{U}}_{1}(h)\times\ldots\times\tilde{\mathcal{U}}_{M}(h)$, where for each $m\in\mathcal{M}$, the individual action set $\tilde{\mathcal{U}}_{m}(h)$ is given by:
@@
Here, for each $m\in\mathcal{M}$, $\delta_{m}\in[0,1]$ and $\zeta_{m}\geq 1$ are chosen conservatively to ensure that the action space is restricted only in clearly suboptimal regions of the state space. In all numerical experiments, we set $\delta_{m}\equiv 0.5$ and $\zeta_{m}\equiv 1.5$, so that PM is prohibited only when $x_{m}$ is well below the OPM threshold, and doing nothing is prohibited only when $x_{m}$ is substantially above the PM threshold.
The feature representation of a state depends on the information available to the manager. We propose a feature design that communicates the state information for each asset in a manner that is customized to the specific asset for which we are currently determining an action. Verleijsdonk et al. [2024, Section 7.1] demonstrate that using such a feature design results in less training variability and consistently produces policies that perform significantly better.
@@
that is, the feature vector $f^{\textbf{L}_{2}}_{1}(h)$ includes an information block $(x_{m},\lambda_{m},p_{m},\iota_{m})$ for each $m\in\mathcal{M}$ that can be derived from $h$, along with one additional feature. Here, $x_{m}$ is the observed degradation level of the $m$-th asset, and the entries $\lambda_{m}$ and $p_{m}$ denote its degradation parameters (cf. Example 1). The last block entry $\iota_{m}$ indicates whether we are currently selecting an action for asset $m$. Lastly, the additional feature $\eta$ indicates whether there is an opportunity for OPM. More specifically, $\eta$ equals $1$ if maintenance on at least one asset is mandatory or if the maintenance action has already been selected in the current decision epoch, and 0 otherwise. Note that the dimension of the feature vector is $r=4M+1$.
@@
$(\lambda^{\textrm{Bayes}}_{m},p^{\textrm{Bayes}}_{m})=\left(\frac{\alpha_{m}(t)}{\beta_{m}(t)},\frac{r_{m}(t)}{r_{m}(t)+s_{m}(t)}\right),$ (9)
where $\alpha_{m}(t),\beta_{m}(t),r_{m}(t)$ and $s_{m}(t)$ are defined in Eqs. (4)–(7).
@@ 5 Simulation study
We compare the performance of each proposed solution approach with that of a policy optimized under information level $\textbf{L}_{2}$. We restrict our analysis to the case where $M=2$ as it facilitates visualization of the trained policies. Furthermore, we assume that all component replacements come from the same pool of components, meaning that $\Lambda_{m}\overset{\textrm{d}}{\equiv}\Lambda$ and $\Phi_{m}\overset{\textrm{d}}{\equiv}\Phi$. For each instance of the simulation study, the true hyperparameters of the gamma distribution $\Lambda$ and the beta distribution $\Phi$ that model the population heterogeneity are denoted by $\alpha$, $\beta$, $r$ and $s$, and are listed in Table 1. The means of the distributions $\Lambda$ and $\Phi$ are denoted by $\mu_{\Lambda}$ and $\mu_{\Phi}$, respectively. These instances are selected to reflect increasing volatility in component heterogeneity, represented by an increasing coefficient of variation (CV) of the distributions, as well as a range of representative cost parameters.
Instance $M$ $\xi_{m}$ $\mu_{\Lambda}$ $\textrm{CV}_{\Lambda}$ $\mu_{\Phi}$ $\textrm{CV}_{\Phi}$ $\alpha$ $\nicefrac{{1}}{{\beta}}$ $r$ $s$ $c_{m}^{\textrm{PM}}$ $c_{m}^{\textrm{CM}}$ $c^{\textrm{ST}}$ $\gamma$ I.1 22 2020 11 $0.3$ $0.5$ $0.01$ $11.11$ $0.09$ $4999.5$ $4999.5$ 1 5 1 0.99 I.2 22 2020 11 $0.6$ $0.5$ $0.02$ $2.78$ $0.36$ $1249.5$ $1249.5$ 1 10 1 0.99
Table 1: Hyperparameter settings and cost structures (adopted from Drent et al. [2023, Section 5]) considered in the simulation study.
All performance results in this section are retrieved using $10^{6}$ repetitions of length $10^{3}$ time units. The reported half-widths represent asymptotic $95\%$ confidence intervals and account solely for variability within the model.